### v1

# Connected Partitions via Connected Dominating Sets

Abstract
The classical theorem due to Győri and Lovász states that any $k$-connected graph $G$ admits a partition into $k$ connected subgraphs, where each subgraph has a prescribed size and contains a prescribed vertex, as soon as the total size of target subgraphs is equal to the size of $G$. However, this result is notoriously evasive in terms of efficient constructions, and it is still unknown whether such a partition can be computed in polynomial time, even for $k=5$.
We make progress towards an efficient constructive version of the Győri–Lovász theorem by considering a natural weakening of the $k$-connectivity requirement. Specifically, we show that the desired connected partition can be found in polynomial time, if $G$ contains $k$ disjoint connected dominating sets. As a consequence of this result, we give several efficient approximate and exact constructive versions of the original Győri–Lovász theorem:
  • On general graphs, a Győri–Lovász partition with $k$ parts can be computed in polynomial time when the input graph has connectivity $\Omega(k\cdot\log^{2}n)$;
  • On convex bipartite graphs, connectivity of $4k$ is sufficient;
  • On biconvex graphs and interval graphs, connectivity of $k$ is sufficient, meaning that our algorithm gives a “true” constructive version of the theorem on these graph classes.
keywords:
Győri–Lovász theorem, connected dominating sets, graph classes
category:
[…]
1 Introduction
Partitioning a graph into connected subgraphs is a problem that naturally arises in various application areas such as image processing, road network decomposition and parallel graph systems for supporting computations on large-scale graphs [32, 33, 19, 23]. Usually in such applications, there is a certain number of parts $k$ that the partition should have, as well as certain constraints on the size of each part. In particular, a prominent objective is for the size of the parts to be as equal as possible and, therefore, for the partition to be as balanced as possible, known as Balanced Connected Partition problem ($BCP_{k}$). This problem is proven to be NP-complete [21] and even hard to approximate [15] and, hence, it has been studied in restricted graph classes such as grid graphs for which it was again proven to be NP-hard [3].
From a theoretical point of view, partitioning a graph into connected components each of a certain, chosen in advance, size is of independent interest as it gives us an insight into the structure of the graph. With the simple example of a star graph, one can observe that it is not possible to partition it into connected sets where more than one of them has size $2$ or more. In general, if there exists a set of $t$ vertices whose removal disconnects the graph (called a separator), and if the number of parts $k$ is greater than $t$, then the $k$ parts of a connected partition might not necessarily be of arbitrary size. Graphs that do not contain a separator of size strictly less than $t$ are called $t$-connected.
The essence of this observation was captured in 1976 by Győri [24] and Lovász [31] who independently proved that the $k$-connectivity of a graph is a necessary and sufficient condition for any partition of a specific size to be possible, even when each part is additionally required to contain a specified vertex called a terminal. This was later generalized to weighted graphs by Chandran et al. [13], Chen et al. [14] and Hoyer [26]. Formally, the result of Győri and Lovász is stated as follows.
Theorem (Győri–Lovász Theorem).
Let $k\geq 1$ be an integer, $G=(V,E)$ be a $k$-connected graph, $c_{1},\ldots,c_{k}\in V$ be distinct vertices of $G$, and $n_{1},\ldots,n_{k}\in\mathbb{N}$ be integers such that $\sum_{i=1}^{k}n_{i}=|V|$. Then $V$ can be partitioned into $k$ disjoint subsets $V_{1},\ldots,V_{k}$ such that $|V_{i}|=n_{i}$, $c_{i}\in V_{i}$, and the induced subgraph $G[V_{i}]$ on $V_{i}$ is connected, for all $i\in[k]$.
Although the proof Győri [24] is constructive, the resulting algorithm runs in exponential time. To this day, it remains open whether finding a Győri–Lovász partition of a given $k$-connected graph could be done in polynomial time in general. On the other hand, a few polynomial algorithms have been developed for special cases. The original research direction is concerned with smaller values of $k$. Specifically, in 1990, Suzuki et al. [38, 39] provided a polynomial algorithm for the cases $k=2$ and $k=3$. Later, in 1993, Wada et al. [40] extended the result for $k=3$, and, in 1997, Nakano et al. [34] showed a linear-time algorithm for the case where $k=4$ and the graph is planar with all the terminals located on the same face of the planar embedding. More recently, in 2016, Hoyer and Thomas [26] presented a polynomial-time algorithm for the case $k=4$ without additional restrictions on the graph structure. It is still open whether a polynomial-time algorithm exists for any fixed value of $k$ starting from $5$. Another research direction focuses on polynomial-time constructions for the Győri–Lovász theorem on specific graph classes, for all values of $k$. Recently, Casel et al. [10] developed a polynomial-time algorithm for chordal graphs and (with a small additive violation of the size constraint) for $\text{HHI}^{2}_{4}$-free graphs. Additionally, Borndörfer et al. [7] have shown an approximate version of the Győri–Lovász theorem on general graphs, where the size of each part is within a factor of $3$ from the given target.
In the light of the challenges above, it is natural to ask whether efficient constructions exist for more restrictive versions of the Győri–Lovász theorem. Censor-Hillel, Ghaffari, and Kuhn [12] pioneered the notion of a connected dominating set (CDS) partition as a proxy for vertex connectivity of the graph. Formally, a graph admits a CDS partition of size $k$, if there exists a partition of the vertices into $k$ sets, so that each set forms a dominating set of the graph, and induces a connected subgraph. It is easy to see that any graph that admits a CDS partition of size $k$, is also $k$-connected. The converse, unfortunately, does not hold; however, Censor-Hillel, Ghaffari, and Kuhn show that there is only a polylogarithmic gap between the connectivity and the size of the maximum CDS partition. Moreover, a suitable CDS partition can be efficiently computed. The bound was later improved by Censor-Hillel et al. [11].
Theorem 1.1 (Corollary 1.6, [11]).
Every $k$-connected $n$-vertex graph has a CDS partition of size $\Omega(\frac{k}{\log^{2}n})$, and such a partition can be computed in polynomial time.
On the other hand, Censor-Hillel, Ghaffari, and Kuhn [12] show that a logarithmic gap between the connectivity and the size of a CDS partition is necessary.
Theorem 1.2 (Theorem 1.3, [12]).
For any sufficiently large $n$, and any $k\geq 1$, there exist $k$-connected $n$-vertex graphs, where the maximum CDS partition size is $O(\frac{k}{\log n})$.
More recently, Draganić and Krivelevich [18] obtained a tighter bound of $(1+o(1))d/\ln d$ for CDS partitions in $d$-regular pseudorandom graphs.
Our contribution
The core result of our work is an efficient version of the Győri–Lovász theorem that is based on the existence of a CDS partition of size $k$ in the given graph, which strengthens the original requirement of $k$-connectivity.
Theorem 1.3.
There is a polynomial-time algorithm that, given a graph $G=(V,E)$, $k$ distinct vertices $c_{1},\ldots,c_{k}\in V$ of $G$, $k$ integers $n_{1},\ldots,n_{k}\in\mathbb{N}$ such that $\sum_{i=1}^{k}n_{i}=|V|$, and a CDS partition of $G$ of size $k$, returns a partition of $V$ into $k$ disjoint subsets $V_{1},\ldots,V_{k}$ such that $|V_{i}|=n_{i}$, $c_{i}\in V_{i}$, and the induced subgraph $G[V_{i}]$ on $V_{i}$ is connected, for all $i\in[k]$.
However, on its own Theorem 1.3 would have a notable limitation, that computing an adequate CDS partition may in itself be an obstacle—as opposed to connectivity, which can always be computed exactly in polynomial time. Fortunately, it turns out that a sufficiently large CDS partition can be found efficiently in a range of scenarios. As the first example, Theorem 1.1 due to Censor-Hillel et al. [11] provides such a construction for general graphs. Combining it with our algorithm, we obtain the following efficient version of the Győri–Lovász theorem, which requires higher connectivity, but puts no additional restrictions on the graph.
Corollary 1.4.
There is a polynomial-time algorithm that, given a $\Omega(\frac{k}{\log^{2}n})$-connected $n$-vertex graph $G=(V,E)$, $k$ distinct vertices $c_{1},\ldots,c_{k}\in V$ of $G$, $k$ integers $n_{1},\ldots,n_{k}\in\mathbb{N}$ such that $\sum_{i=1}^{k}n_{i}=|V|$, returns a partition of $V$ into $k$ disjoint subsets $V_{1},\ldots,V_{k}$ such that $|V_{i}|=n_{i}$, $c_{i}\in V_{i}$, and the induced subgraph $G[V_{i}]$ on $V_{i}$ is connected, for all $i\in[k]$.
Furthermore, we investigate the cases where Theorem 1.3 can be applied to obtain a tighter connectivity gap. Following the results of Censor-Hillel, Ghaffari, and Kuhn [12], the case of general graphs cannot be improved beyond a logarithmic factor, as there exist $k$-connected graphs that do not admit CDS partitions of size $O(k\cdot\log n)$, see Theorem 1.1. Therefore, putting additional restrictions on the input graph is necessary.
The first natural target here would be efficient constructions for the Győri–Lovász theorem on certain graph classes in its exact version, that is, requiring connectivity exactly $k$. In the light of Theorem 1.3, it is sufficient to show a statement of the following form, that any $k$-connected graph in the class $\mathcal{G}$ admits a CDS partition of size $k$. This would imply that the Győri–Lovász theorem admits an efficient construction on graphs in $\mathcal{G}$. We identify two such cases, biconvex graphs and interval graphs.
Theorem 1.5.
Let $G$ be a $k$-connected biconvex graph. A CDS partition of $G$ of size $k$ exists, and can be computed in polynomial time.
Theorem 1.6.
Let $G$ be a $k$-connected interval graph. A CDS partition of $G$ of size $k$ exists, and can be computed in polynomial time.
Coupled with the algorithm of Theorem 1.3, this gives an efficient construction for the Győri–Lovász theorem on biconvex graphs and interval graphs.
Corollary 1.7.
There is a polynomial-time algorithm that, given a $k$-connected biconvex graph $G=(V,E)$, $k$ distinct vertices $c_{1},\ldots,c_{k}\in V$ of $G$, $k$ integers $n_{1},\ldots,n_{k}\in\mathbb{N}$ such that $\sum_{i=1}^{k}n_{i}=|V|$, returns a partition of $V$ into $k$ disjoint subsets $V_{1},\ldots,V_{k}$ such that $|V_{i}|=n_{i}$, $c_{i}\in V_{i}$, and the induced subgraph $G[V_{i}]$ on $V_{i}$ is connected, for all $i\in[k]$.
We note that the efficient construction for chordal graphs was identified by Casel et al. [10], implying the same result for interval graphs, since interval graphs form a subclass of chordal graphs. Nonetheless, we state the algorithmic consequence of Theorem 1.6 next for completeness.
Corollary 1.8.
There is a polynomial-time algorithm that, given a $k$-connected interval graph $G=(V,E)$, $k$ distinct vertices $c_{1},\ldots,c_{k}\in V$ of $G$, $k$ integers $n_{1},\ldots,n_{k}\in\mathbb{N}$ such that $\sum_{i=1}^{k}n_{i}=|V|$, returns a partition of $V$ into $k$ disjoint subsets $V_{1},\ldots,V_{k}$ such that $|V_{i}|=n_{i}$, $c_{i}\in V_{i}$, and the induced subgraph $G[V_{i}]$ on $V_{i}$ is connected, for all $i\in[k]$.
Unfortunately, there is a limit to the results of the form above, as already graph classes such as chordal graphs and convex bipartite graphs do not possess the property that $k$-connectivity implies the existence of a CDS partition of size $k$ (see Figure 1 for examples).
Figure 1: Examples of a $2$-connected convex bipartite graph and a $2$-connected chordal graph that do not admit a CDS partition of size $2$
On the positive side, we are able to retain the productivity of Theorem 1.3 even in such cases. We focus on the class of convex bipartite graphs and show that $4k$-connected convex bipartite graphs always admit a CDS partition of size $k$.
Theorem 1.9.
Let $G$ be a $4k$-connected convex bipartite graph. A CDS partition of $G$ of size $k$ exists, and can be computed in polynomial time.
As before, combined with the algorithm of Theorem 1.3, this gives us a “$4$-approximate” (in terms of connectivity) efficient version of the Győri–Lovász theorem on this graph class, considerably improving the connectivity requirement of the general case.
Corollary 1.10.
There is a polynomial-time algorithm that, given a $4k$-connected convex bipartite graph $G=(V,E)$, $k$ distinct vertices $c_{1},\ldots,c_{k}\in V$ of $G$, $k$ integers $n_{1},\ldots,n_{k}\in\mathbb{N}$ such that $\sum_{i=1}^{k}n_{i}=|V|$, returns a partition of $V$ into $k$ disjoint subsets $V_{1},\ldots,V_{k}$ such that $|V_{i}|=n_{i}$, $c_{i}\in V_{i}$, and the induced subgraph $G[V_{i}]$ on $V_{i}$ is connected, for all $i\in[k]$.
Related work
The problem of computing the largest CDS partition of a graph was first introduced by Hedetniemi and Laskar [25], where the maximum size of a CDS partition is referred to as the connected domatic number of the graph. A number of bounds on the size of a CDS partition are due to Zelinka [42], including the fact that it is upper-bounded by the connectivity of the graph. Connected dominating sets have many practical applications for tasks such as routing and load-balancing in networks, for a detailed overview we refer to a survey due to Yu et al. [41]. Connections between CDS partitions and graph rigidity were explored by Krivelevich, Lew and Michaeli [28].
Chordal graphs are graphs that do not admit induced cycles of length more than 3. The class of chordal graphs is well-studied, and is known to possess many useful properties; in particular, many problems that are NP-hard on general graphs turn out to be polynomial-time solvable on chordal graphs, such as computing chromatic number, independence number, treewidth [35]. Interval graphs are intersection graphs of intervals on a real line, and form a subclass of chordal graphs [29].
Convex bipartite graphs [^1] Often referred to simply as convex graphs in the literature. were introduced in 1967 by Glover [22] as they naturally arise in various industrial and scheduling applications. Formally, a bipartite graph $G=(A\cup B,E)$ is called convex when there exists an ordering $\sigma$ of the vertices in $A$ such that the neighborhood of any vertex of $B$ is consecutive in $A$ under $\sigma$. Restricting to the class of convex bipartite graphs has been proven fruitful for many problems such as maximum independent set, minimum feedback set, (dynamic) maximum matching and independent dominating set [36, 9, 30, 27, 37, 16]. On the other hand, several natural problems are known to be NP-complete on convex bipartite graphs. Most relevant to our work is the example of the $k$-path partition problem, showed to be NP-complete on convex bipartite graphs by Asdre and Nikolopoulos [2]. Belmonte and Vatshelle [4] show that convex bipartite graphs have bounded mim-width, which provides a unified perspective on polynomial-time algorithms for various classes. Bonomo-Braberman et al. [5] generalize this perspective further to extended notions of convexity.
Bipartite graphs with an ordering satisfying the convexity condition for both parts, are called biconvex. Biconvex graphs are thus a subclass of convex bipartite graphs. Abbas and Stewart studied properties of vertex orderings in biconvex graphs, yielding in particular that the vertex ranking problem is polynomial-time solvable on biconvex graphs [1]. Apart from their structural and algorithmic interest, biconvex graphs and convex bipartite graphs have been proven useful in practice, as they model well situations such as task allocation problems [22, 27]. In terms of other well-studied graph classes, convex bipartite graphs and biconvex graphs form a subclass of chordal bipartite graphs, which have a natural analogue of the chordal property among the bipartite graphs. For a more detailed introduction to the graph classes, their properties and relations between the classes, we refer to a survey by Brandstädt, Bang Le and Spinrad [8].
Paper organization.
In Section 2, we introduce preliminaries that we use throughout the paper. Section 3 is dedicated to the algorithm constructing a Győri–Lovász partition on the basis of a CDS partition. In Sections 4 and 5 we consider biconvex graphs and interval graphs, respectively. We conclude with an overview and selected open problems in Section 6.
2 Preliminaries
All graphs mentioned in this paper are finite, undirected and simple. We use standard graph theory notation, for reference see the book of Diestel [17]. A graph $G=(V,E)$ is defined by its set of vertices $V$, also denoted by $V(G)$, and its set of edges $E$, also denoted by $E(G)$. We say that $|V|$ is the order of $G$, and we usually denote it by $n$. Given a graph $G=(V,E)$, we denote by $P(V)$ the powerset of the vertices of $G$. Given a vertex set $X\in V$ we denote by $N_{G}(X)=\{u\in V\setminus X:uv\in E\text{ and }v\in X\}$ the open neighbourhood of $X$ in $G$ and by $N_{G}[X]=X\cup N(X)$ the closed neighbourhood of $X$ in $G$. If the graph we refer to, is clear by the context we omit $G$ as a subscript. We also refer to $|N(v)|$ as the degree of the vertex $v$, denoted $\deg(v)$. Moreover, given a vertex set $X\in V(G)$ we denote by $G[X]$ the subgraph of $G$ induced by $X$, the graph with vertex set $X$ and with edge set the subset of $E$ that contains all the edges with both endpoints in $X$.
We say that a graph is $k$-connected if the removal of any vertex set of size at most $k-1$ does not disconnect the graph. The maximum value of $k$ for which a graph $G$ is $k$-connected, is called the connectivity of $G$ and is denoted by $\kappa(G)$. We recall the statement of the well-known theorem by Menger next, and refer to the book by Diestel [17] for this version of the theorem and its proof. Note that the proof of Theorem 2.1 is constructive, meaning that the respective collection of paths can be computed in polynomial time [17].
Theorem 2.1 (Menger’s Theorem; Theorem 3.3.6, [17]).
A graph is $k$-connected if and only if it contains $k$ independent paths between any two vertices.
We call a graph $G=(V,E)$ a path if $V=\{v_{1},\ldots,v_{n}\}$ and $E=\{v_{i}v_{i+1}|\text{for }i\in[n-1]\}$. We usually denote a path by $P$ and we call its order, the length of $P$. We call a graph $G=(V,E)$ a cycle if $V=\{v_{1},\ldots,v_{n}\}$ and $E=\{v_{i}v_{i+1},v_{1}v_{n}|\text{for }i\in[n-1]\}$. We call a graph $G=(V,E)$ a tree if it does not contain any cycle as a subgraph. Throughout this paper a vertex set of a tree is usually denoted by $T$, while its edge set is denoted by $E(T)$. Given two graphs $H$, $G$ we say that $G$ is $H$-free if $G$ does not contain $H$ as an induced subgraph. Given two graphs $G$ and $T$, we say that $T$ is a spanning tree of $G$ if $V(G)=V(T)$ and $G[V(T)]$ induces a tree.
Definition 2.2 (Path decomposition).
A path decomposition of a graph $G=(V,E)$ is a sequence of subsets $B_{1},\dots,B_{\ell}$ of $V$ (also called bags) satisfying the following conditions:
  1. 1.
    $\bigcup_{i=1}^{\ell}B_{i}=V$.
  2. 2.
    For every edge $uv\in E$, there exists an index $i$ such that $u,v\in B_{i}$.
  3. 3.
    For every vertex $v\in V$, the set of indices $\{i\mid v\in B_{i}\}$ forms a contiguous subsequence of $\{1,\dots,\ell\}$.
The width of a path decomposition is $\max_{1\leq i\leq\ell}|B_{i}|-1$, and the pathwidth of $G$ is the minimum width over all possible path decompositions of $G$.
A graph is said to be bipartite if its vertices can be partitioned into two sets $A$ and $B$ such that each edge in $G$ has its endpoints in different parts and we denote it as $G=(A\cup B,E)$. A bipartite graph $G=(A\cup B,E)$ is called biconvex bipartite if there exist orderings $\sigma_{A}:A\rightarrow[|A|]$ and $\sigma_{B}:B\rightarrow[|B|]$ such that for each vertex $v\in A\cup B$, the vertices of $N(v)$ are consecutive under the corresponding ordering. These orderings together is called a biconvex ordering of $A\cup B$. A bipartite graph $G=(A\cup B,E)$ is called convex bipartite if there exists an ordering $\sigma:A\rightarrow[|A|]$ such that for every vertex $v\in B$, the vertices of $N(v)$ are consecutive under $\sigma$. Such an ordering $\sigma$ is called convex ordering of $A$ and can be computed in linear time [6]. For biconvex and convex bipartite graphs, given two vertices $u,v$ we denote by $u<v$ the fact that the ordering value of $u$ is less than the one of $v$ (similarly for $u\leq v$). A superclass of convex bipartite graphs are chordal bipartite graphs. We call a bipartite graph $G=(A\cup B,E)$, chordal bipartite, if it is $C_{\ell}$-free for $\ell\geq 6$. A graph $G=(V,E)$ is called interval if there is a set of intervals in the real line such that each vertex denotes an interval and two vertices are adjacent if and only if two intervals overlap.
Given two vertex sets $A,B\in V(G)$ we say that $A$ is dominating to $B$ if every vertex in $B$ is adjacent to some vertex in $A$. Also, when we say that a vertex set is dominating to some graph $G$, we mean that it is dominating to its vertex set $V(G)$. A vertex set $A$ is called a connected dominating set (CDS) if and only if it is a dominating set and $G[A]$ is connected. Given a graph $G$, a collection of vertex sets $\mathcal{V}=\{V_{1},\ldots,V_{l}\}$ such that $V_{i}\subseteq V(G)$ is called a connected dominating set partition (CDS partition) if and only if the sets in $\mathcal{V}$ are pairwise disjoint and $\cup_{U\in\mathcal{V}}U=V(G)$, and each of them is a connected dominating set of $G$. We call $|\mathcal{V}|$ the size of the CDS partition $\mathcal{V}$. Throughout the paper, we also use an equivalent notion of vertex-disjoint dominating trees. Given a CDS partition of a graph it is immediate to get a set of vertex-disjoint dominating trees of the same size by considering the spanning tree of each set in the CDS partition. Moreover, given $k$ disjoint dominating trees, one may obtain a CDS partition of size $k$ by simply adding the remaining vertices arbitrarily to those trees.
3 From CDS Partition to Győri–Lovász Partition
In this section, we develop a polynomial-time algorithm realizing Theorem 1.3 (Algorithm 7). We first describe the intuition behind the main steps of the algorithm, then provide a more detailed description of these steps, and finally define a formal procedure for each step and show its correctness.
Intuition of Algorithm 7: The main challenge of constructing connected partitions from the dominating trees, lies in the distribution of terminal vertices. In particular, if each terminal is contained in a distinct dominating tree, then it suffices to add the vertices of each tree to the corresponding set (until the size demand is met) to create $k$ dominating connected sets and then add the remaining vertices arbitrarily until all the size demands are met, since the connectivity of each set is ensured due to the dominating property of each set.
The difficult case is when a dominating tree contains more than one terminal. Suppose there are $l$ terminals in some dominating tree $T$. First, we assign all the vertices of the tree $T$ to one of the $l$ vertex sets while maintaining the connectivity of each set. There exist $l-1$ dominating trees that have no terminal vertex. Now, we map these $l$ vertex sets to one of these $l-1$ dominating trees. We aim to fulfill the vertex set to its demand while using the corresponding dominating tree.
We use an iterative procedure that first assigns (but does not add) neighbors for each vertex set to these $l-1$ dominating trees, and then we add assigned vertices to each vertex set whose demand is not met by the assignment. Now, to fulfill the demand of each such vertex set, we use its corresponding dominating tree while preserving the connectivity of the vertex set. At the end of this procedure, each part is either full or contains a whole dominating tree (apart from the one that was not assigned to any tree). The remaining vertices are arbitrarily added to the non-full sets, as each contains a dominating tree, apart from the one that has not been assigned a tree, for which we have ensured that there are enough adjacent vertices left.
Description of Algorithm 7: The Algorithm receives as input a $k$-connected graph $G$, a set of $k$, vertex disjoint, dominating trees $\mathcal{T}=\{T_{1},\ldots,T_{k}\}$, a set of $k$ distinct terminal vertices $C=\{c_{1},\ldots,c_{k}\}$ and $k$ natural numbers $\mathcal{N}=\{n_{1},\ldots,n_{k}\}$ such that $\sum_{i=1}^{k}n_{i}=|V(G)|$. Moreover, we assume without loss of generality that $n_{i}\geq 2$ for all $i\in[k]$ as otherwise, if there is a set with size demand $1$, we simply output the corresponding terminal and run the algorithm for the remaining $(k-1)$-connected graph. It is important to note that during this description of the algorithm, every time we add a vertex to a vertex set, it is implied that we also check whether if the condition of having $l$ full sets say $V_{1},\ldots,V_{l}$, intersecting exactly $l$ of the dominating trees, say $T_{1},\ldots,T_{l}$ is met. If yes, then we output those $l$ full sets and restart the algorithm for $G-\cup_{i=1}^{l}V_{i}$ and $k-l$, $\mathcal{T}=\mathcal{T}\setminus\{T_{1},\ldots,T_{l}\}$, $C=C\setminus\{c_{1},\ldots,c_{l}\}$ and $\mathcal{N}=\mathcal{N}\setminus\{n_{1},\ldots,n_{l}\}$. For simplicity, we do not include this in every step of this description; it is, however, included in the pseudocode corresponding to each step.
  1. 1.
    The trees of $\mathcal{T}$ are categorized in the sets $\mathcal{T}_{0}$, $\mathcal{T}_{1}$ and $\mathcal{T}_{>1}$ depending on whether a tree contains 00, $1$ or more than $1$ terminal vertices respectively. Moreover, if some terminal does not belong to any tree, we add it to one in $\mathcal{T}_{0}$ and move that tree from $\mathcal{T}_{0}$ to $\mathcal{T}_{1}$. We set each vertex set $V_{i}$ to contain the terminal vertex $c_{i}$ for all $i\in[k]$ (Algorithm 1).
  2. 2.
    For each tree in $\mathcal{T}_{1}\cup\mathcal{T}_{>1}$, we add its vertices one by one to a vertex set that it is adjacent to (based only on the adjacencies occurring by the tree considered). Next, we add the vertices of $V(G)\setminus V(\mathcal{T})$ to an adjacent set one by one (Algorithm 2).
  3. 3.
    For each tree $T\in\mathcal{T}_{>1}$ that contains $l$ terminals, we create a tree set $\mathcal{T}^{\prime}$ containing $T$ and $l-1$ trees from $\mathcal{T}_{0}$, which we then remove from $\mathcal{T}_{>1}$ and $\mathcal{T}_{0}$ respectively. Among those sets, we chose one that has enough vertices available to satisfy the size demands of all of the $l$ vertex sets intersecting $T$. (Algorithm 3)
  4. 4.
    Each vertex of $\mathcal{T}^{\prime}\setminus T$ is assigned through a function $Vlabel$ to an adjacent vertex set among the $l$ intersecting $T$. Depending on whether the amount of vertices assigned to some vertex set is less than the vertices needed to be added for the set to satisfy its demand or not, the set is categorized in the set $Under$ or $Over$ respectively (Algorithm 3).
  5. 5.
    Through the function $Tlabel$ each set in $Under$ is assigned one of the trees in $\mathcal{T}^{\prime}\setminus T$. Then, one vertex of each such tree adjacent to the corresponding set is added. Priority is given to the vertices assigned to this set through the function $Vlabel$. However, if no such vertex exists, another adjacent vertex is arbitrarily chosen to be added, and we change the value of this vertex in the function $Vlabel$ accordingly. If this results in a set moving from $Over$ to $Under$, we again assign a tree to this set (that has not been assigned to any other set) though the function $Tlabel$ (Algorithm 4).
  6. 6.
    For each tree assigned to some set, we consider its vertices not belonging in that set one by one and add them to that set as long as the set connectivity is preserved. Again, if, during this procedure, a set moves from $Over$ to $Under$, go to the previous Step. Repeat this procedure until each set in $Under$ is either full or contains a whole tree (Algorithm 5).
  7. 7.
    Add each of the remaining vertices, the ones not added to any set, to an adjacent full set one by one (Algorithm 6).
  8. 8.
    Output the set of $l$ full connected sets that intersect at most $l$ dominating trees and restart the algorithm for the updated graph $G$ and respective values.
Algorithm 1 Categorize Trees (Step 1)
1:A set of $k$ dominating trees $\mathcal{T}=\{T_{1},\ldots,T_{k}\}$, a set of $k$ terminals $C=\{c_{1},\ldots,c_{k}\}$
2:for $i\leftarrow 1\textrm{ to }k$ do
3:     if $|T_{i}\cap C|>1$ then
4:         $\mathcal{T}_{>1}\leftarrow\mathcal{T}_{>1}\cup T_{i}$
5:     else if $|T_{i}\cap C|=1$ then
6:         $\mathcal{T}_{1}\leftarrow\mathcal{T}_{1}\cup T_{i}$
7:     else
8:         $\mathcal{T}_{0}\leftarrow\mathcal{T}_{0}\cup T_{i}$
9:for $i\leftarrow 1\textrm{ to }k$ do
10:     $V_{i}\leftarrow c_{i}$
11:     if $c_{i}\not\in V(\mathcal{T})$ then
12:         Let $T\in\mathcal{T}_{0}$. Then $T\leftarrow T\cup c_{i}$, $\mathcal{T}_{0}\leftarrow\mathcal{T}_{0}\setminus T$ and $\mathcal{T}_{1}\leftarrow\mathcal{T}_{1}\cup T$      return $\{\mathcal{T}_{0}$, $\mathcal{T}_{1}$, $\mathcal{T}_{>1}\}$, $\{V_{1},\ldots,V_{k}\}$
Lemma 3.1.
Given $k$ dominating trees, and a set of $k$ distinct terminal vertices as an input, Algorithm 1 (Step 1) returns $k$ dominating trees such that each terminal vertex is part of one of those trees and $k$ vertex sets, each containing a distinct terminal vertex.
Proof 3.2.
Notice that up to line 9 terminal vertices that are part of some tree define in which of the three tree categories $\mathcal{T}_{0}$, $\mathcal{T}_{1}$, $\mathcal{T}_{>1}$ each tree belongs in (containing 00, $1$ and more than $1$ terminal vertices respectively). Since each tree is dominating, each terminal is adjacent to each tree. Hence, by the loop at line 9, each terminal vertex is added to a tree that does not contain any terminal so far, and due to their adjacency, the resulting graph is still connected, and hence there is the corresponding spanning tree which is dominating to the rest of the graph. It also reassigns the tree to a suitable category, denoting its current number of terminals. Lastly, notice that each terminal vertex is added to the desired vertex set in the corresponding loop of the algorithm.
Algorithm 2 Add Vertices (Step 2)
1:A set of $k$ trees $\mathcal{T}=\{T_{1},\ldots,T_{k}\}$ separated in three categories, $\{\mathcal{T}_{0},\mathcal{T}_{1},\mathcal{T}_{>1}\}$, a set of $k$ terminals $C=\{c_{1},\ldots,c_{k}\}$, $k$ connected vertex sets $\{V_{1},\ldots,V_{k}\}$ such that $c_{i}\in V_{i}$ for all $i\in[k]$, a set of $k$ natural numbers $N=\{n_{1},\ldots,n_{k}\}$ and a set of vertices $U$
2:for $i\leftarrow 1\textrm{ to }k$ $\triangleright$ Add Vertices of Trees
3:     if $T_{i}\in\mathcal{T}_{1}\cup\mathcal{T}_{>1}$ then
4:         Denote by $V_{1},\ldots,V_{l}$ the $l$ sets that intersect $T_{i}$, and by $U_{T_{i}}$ the open neighborhood of these sets on $T_{i}$, formally $N_{T_{i}}(\cup_{j=1}^{l}V_{j})$
5:         while $U_{T_{i}}\neq\emptyset$ do
6:              Let $v\in U_{T_{i}}$ and $V_{s}\in\{V_{1},\ldots,V_{l}\}$ such that $v\in N_{T_{i}}(V_{s})$.
7:              $V_{s}\leftarrow V_{s}\cup v$, $U_{T_{i}}\leftarrow N_{T_{i}}(\cup_{j=1}^{l}V_{j})$
8:              if $|V_{s}|=n_{s}$ then
9:                  Output $V_{s}$
10:                  Run Algorithm 7 with input $(G\setminus V_{s},C\setminus c_{s},N\setminus n_{s},\mathcal{T}\setminus T_{i})$
11:while $U\neq\emptyset$ $\triangleright$ Add Non-Tree Vertices
12:     Let $v\in U$ and $V_{s}\in\{V_{1},\ldots,V_{k}\}$ such that $v\in N(V_{s})$
13:     $V_{s}\leftarrow V_{s}\cup v$, $U\leftarrow U\setminus v$
14:     if $|V_{s}|=n_{s}$ then
15:         Output $V_{s}$
16:         Run Algorithm 7 with input $(G\setminus V_{s},C\setminus c_{s},N\setminus n_{s},\mathcal{T}\setminus T_{i})$ where $T_{i}$ is the dominating tree in $\mathcal{T}$ such that $c_{s}\in T_{i}$.      return $\{V_{1},\ldots,V_{k}\}$
Lemma 3.3.
Given $k$ dominating trees of a graph $G$ separated in three categories $\mathcal{T}_{0},\mathcal{T}_{1},\mathcal{T}_{>1}$, and $k$ distinct connected vertex sets, each containing one terminal vertex, Algorithm 2 adds each vertex of the trees in $\mathcal{T}_{1}\cup\mathcal{T}_{>1}$ to some vertex set such that each vertex set induces a connected graph. When this algorithm runs as part of Algorithm 7 (Step 2), then either each vertex of each tree in $\mathcal{T}_{1}\cup\mathcal{T}_{>1}$ is added to some vertex set inducing a connected graph or the algorithm restarts for a reduced value of $k$.
Proof 3.4.
Notice that when running Algorithm 2 as part of Algorithm 7, the loop in line 2 considers each tree that contains at least one terminal separately. Then, in the while loop, each vertex of the tree that is adjacent to a non-full set is added to that vertex set, hence preserving the connectivity of each set formed. This procedure terminates when all the vertices of the considered tree are added to some vertex set, and each of those sets is not full or because one of those met its size demand.
If it terminates due to the second condition, meaning some set $V$ becomes full, notice that $V$ is connected, and its removal reduces the connectivity of the graph by at most one because it does not contain any vertex from $k-1$ dominating trees, and hence, those trees cause the graph $G-V$ to be $(k-1)$-connected.
Lemma 3.5.
Given $k$ connected vertex sets $\{V_{1},\ldots,V_{k}\}$ such that their union induces a collection of connected dominating to $U$ sets and $k$ natural numbers $\{n_{1},\ldots,n_{k}\}$, Algorithm 2 adds each vertex of $U$ to some neighboring set $V_{i}\in\{V_{1},\ldots,V_{k}\}$. When this algorithm runs as part of Algorithm 7 (Step 2), then either each vertex of each $U$ is added to some set inducing a connected graph or the algorithm restarts for a reduced value of $k$.
Proof 3.6.
Notice that the loop in line 11 considers every vertex in $U$. Any such vertex is adjacent to some vertex set because, due to the input, their union induces a dominating to $U$ set. Note that when Algorithm 2 runs as part of Algorithm 7, those sets indeed induce sets that are dominating to $U$, as they are part of dominating to $G$ trees. Hence, due to Lemma 3.1 and 3.3, there is at least one dominating tree whose vertices are all added to those connected sets. Each vertex of $U$ is added to a set, increasing its size by one and preserving its connectivity. This loop terminates when all vertices of $U$ are added to some connected vertex set or if one set meets its size demand. In the second case Algorithm 7 restarts for a reduced value of $k$.
Note that if the loop terminates due to the second condition when run as part of Algorithm 7, the full set causing this termination, is connected. Also, its removal reduces the connectivity of the graph by at most one because it does not contain any vertex from $k-1$ dominating trees, and hence those trees make the graph $(G\setminus V_{s})$ to be $(k-1)$-connected.
Lemma 3.7.
Let $G$ be a graph, $n_{1},\ldots,n_{k}$ be $k$ natural numbers such that $\sum_{i=1}^{k}n_{i}=|V(G)|$, $k$ dominating trees partitioned into three sets $\mathcal{T}_{0},\mathcal{T}_{1},\mathcal{T}_{>1}$, and $V_{i}$ be a connected vertex set containing terminal $c_{i}$, where $i\in[k]$, are given as an input to Algorithm 3 (Step 3). Then, there exists a tree $T\in\mathcal{T}_{>1}$ with terminals $c_{1},c_{2},\ldots,c_{l}$ and $l-1$ trees $T^{\prime}_{1},\ldots,T^{\prime}_{l-1}\in\mathcal{T}_{0}$ such that $|T|+\sum_{i=1}^{l-1}|T_{i}^{\prime}|\geq\sum_{i=1}^{l}n_{i}-|V_{i}|$. Moreover, $G[V_{i}\cup Vlabel(V_{i})]$ is a connected graph for each $i\in[l]$ and there is at least one set in $Over$.
Proof 3.8.
Notice that when Algorithm 3 runs as a part of Algorithm 7, it essentially groups each tree containing more than one terminal with the respective number of trees containing no terminals such that the tree set contains an equal number of terminal vertices to dominating trees. We do this by removing $l-1$ trees from $\mathcal{T}_{0}$ any time a tree $T\in\mathcal{T}_{>1}$ is considered, where $l$ is the number of terminal vertices in $T$. We know that enough such trees exist because we have a total of $k$ trees and $k$ terminals.
Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{s}$ be those sets of trees (the ones originally in $\mathcal{T}_{0}$) and assume for a contradiction that none of those sets satisfies the condition of line 8. Moreover, due to Lemma 3.5, we know that all the vertices not originally in $\mathcal{T}$ and the ones in some tree in $\mathcal{T}_{1}\cup\mathcal{T}_{>1}$ are contained in some non-full set. Denote the set of vertex sets created so far, meaning $\{V_{1},\ldots,V_{k}\}$ by $\mathcal{V}$ and by $n_{\mathcal{V}}$ the sum of their size demands. Hence, $|V(G)|=\sum_{i=1}^{s}\sum_{P\in\mathcal{P}_{i}}|V(P)|+\sum_{H\in\mathcal{V}}|V(H)|$. Due to the previous observation and our assumption $|V(G)|<|V(G)|-n_{\mathcal{V}}+n_{\mathcal{V}}$, which leads us to the desired contradiction.
Moreover, because a vertex $u$ is added to a set $Vlabel(V)$ only if it is adjacent to $V$ based only on the adjacencies of the vertices of $V$ that also belong in the tree $T$ and each of the sets $V$ is connected, the graph $G[V_{i}\cup Vlabel(V_{i})]$ is also a connected graph.
Next, we want to show that there exists a set in $Over$. Assume for a contradiction that there is no set in $Over$ and hence $n_{i}-|V_{i}|>|Vlabel(V_{i})|$ for all $i\in[l]$. It follows $\sum_{i=1}^{l}n_{i}-\sum_{i=1}^{l}|V_{i}|>\sum_{i=1}^{l}|Vlabel(V_{i})|\Rightarrow|T|+\sum_{i=1}^{l-1}|T_{i}^{\prime}|-\sum_{i=1}^{l}|V_{i}|>\sum_{i=1}^{l-1}|T_{i}^{\prime}|\Rightarrow|T|>\sum_{i=1}^{l}|V_{i}|$, which is a contradiction as $T\subseteq\cup_{i=1}^{l}V_{i}$ by Lemma 2.
Algorithm 3 Choose a Tree Set (Steps 34)
1:A graph $G$, three sets of dominating trees $\mathcal{T}_{0},\mathcal{T}_{1},\mathcal{T}_{>1}$, a set of $k$ vertices $C=\{c_{1},\ldots,c_{k}\}$, a set of $k$ natural numbers $N=\{n_{1},\ldots,n_{k}\}$ and $k$ connected vertex sets $\{V_{1},\ldots,V_{k}\}$ such that $c_{i}\in V_{i}$ for all $i\in[k]$
2:Initialize function $Vlabel:\{V_{1},\ldots,V_{k}\}\rightarrow P(V(G))$ to be the empty set for each set $V_{i}$, $i\in[k]$
3:while $\mathcal{T}_{0}\neq\emptyset$ $\triangleright$ Step 3
4:     Let $T\in\mathcal{T}_{>1}$ that contains $l$ terminals $\{c_{1},\ldots,c_{l}\}\in C$ and let $V_{1},\ldots,V_{l}$ contains $c_{1},\ldots,c_{l}$, respectively.
5:     Pick $T_{1}^{\prime},\ldots,T_{l-1}^{\prime}\in\mathcal{T}_{0}$.
6:     $\mathcal{T}_{>1}\leftarrow\mathcal{T}_{>1}\setminus T$
7:     $\mathcal{T}_{0}\leftarrow\mathcal{T}_{0}\setminus\{T_{1}^{\prime},\ldots,T_{l-1}^{\prime}\}$.
8:     if $|T|+\sum_{i=1}^{l-1}|T_{i}^{\prime}|\geq\sum_{i=1}^{l}n_{i}-|V_{i}|$ then
9:         for $i\leftarrow 1\textrm{ to }l-1$ do
10:              for every vertex $v\in T_{i}^{\prime}$ $\triangleright$ Step 4
11:                  Let $V_{i}\in\{V_{1},\ldots,V_{k}\}$ such that $v\in N_{G}(V_{i}\cap T)$
12:                  $Vlabel(V_{i})\leftarrow Vlabel(V_{i})\cup v$
13:         $Over\leftarrow\emptyset$, $Under\leftarrow\emptyset$
14:         for $i\leftarrow 1\textrm{ to }l$ do
15:              if $n_{i}-|V_{i}|\leq|Vlabel(V_{i})|$ then
16:                  $Over\leftarrow Over\cup V_{i}$
17:              else
18:                  $Under\leftarrow Under\cup V_{i}$                             return $\{T,T^{\prime}_{1},\ldots,T^{\prime}_{l-1}\},Over,Under,Vlabel$
Notice that in the initial loop of Algorithm 4, each vertex assigned through the function $Vlabel$ to some set in $Under$ is added to the corresponding set. Here, even though we add vertices to a set, it is not necessary to check whether that set meets its size demand, and as a result, we need to restart the algorithm due to the way the sets in $Under$ are defined. Moreover, since the assignment $Vlabel$ is based on vertices of trees being adjacent to the part of the vertex set that intersects $T$, the vertex sets that occur from this addition are connected. Lastly, each vertex set is assigned a distinct tree in this loop through the function $Tlabel$. We know that enough available trees exist for this assignment due to Lemma 3.7.
Lemma 3.9.
Let a graph $G$, $k$ natural numbers $N=\{n_{1},\ldots,n_{k}\}$ such that $\sum_{i=1}^{k}n_{i}=|V(G)|$, a set of $l$ vertices $C=\{c_{1},\ldots,c_{l}\}$, $l$ connected vertex sets $\{V_{1},\ldots,V_{l}\}$, such that $c_{i}\in V_{i}$ for all $i\in[l]$, a tree $T$ dominating to $G$ and $l-1$ dominating trees $\{T_{1}^{\prime},\ldots,T_{l-1}^{\prime}\}$ and a function $Vlabel:\{V_{1},\ldots,V_{l}\}\to P(V(G))$, are given as input to Algorithm 4. Then, after the loop in line 8 (Step 5) each vertex set is either in $Over$ or is assigned to a tree through the function $Tlabel$ and also contains a vertex from that tree.
Proof 3.10.
In the loop in line 8, we consider each tree assigned to some vertex set and first check whether this vertex set already contains some vertex of the tree. If yes, our lemma is satisfied. Otherwise, since the tree is dominating, there exists at least one vertex adjacent to the part of the vertex set intersecting $T$. This vertex is already added to some vertex set in $Under$ or assigned to one in $Over$. If it belongs in some set in $Over$, then we simply add it to the vertex set. Then we remove this vertex from the set that occurs from the assignment $Vlabel$ for the set previously in $Over$, and we check to see if it now meets the condition for this set belonging to $Under$. If yes, we go to line 3 to deal with this set similarly. Repeating this procedure for the other sets in $Under$ does not affect them as every vertex set currently in $Under$ already contains all the vertices assigned to it through the function $Vlabel$. Moreover, since in this step, we add a vertex to a vertex set, we also check if this set becomes full and intersects exactly one dominating tree. If that is true, we restart the algorithm for the adjusted values of $k$, $G$, $\mathcal{T}$, and $\mathcal{C}$.
Otherwise, if the adjacent vertex is part of some other set in $Under$, we add it to our current set and remove it from the previous one (again checking if the condition for restarting the algorithm is met). Moreover, we remove it from the set created through the function $Vlabel$ to avoid this vertex from existing in two sets (in case we go to line 3) in some other iteration of this loop. Note that sets initially in $Under$ might be added in $Full$ or removed from it depending on whether we add or remove vertices from them, but they will never go from $Under$ to $Over$ as we need to maintain the property that each set in $Over$ contains no vertices from $T^{\prime}_{1},\ldots,T^{\prime}_{l-1}$.
Algorithm 4 Assign Trees (Step 5)
1:A $k$-connected graph $G$, a set of $l$ dominating trees $\{T,T^{\prime}_{1},\ldots,T^{\prime}_{l-1}\}$, a set of $l$ vertices $C=\{c_{1},\ldots,c_{l}\}$, $l$ connected vertex sets $\{V_{1},\ldots,V_{\}}$, such that $c_{i}\in V_{i}$ for all $i\in[l]$, separated in two sets $Under$, $Over$, a function $Vlabel:\{V_{1},\ldots,V_{k}\}\to P(V(G))$, a set of connected vertex sets $Full$ and a set of $l$ natural numbers $N=\{n_{1},\ldots,n_{l}\}$
2:Initialize the value of the function $Tlabel:\{V_{1}\ldots,V_{l}\}\to\{T_{1},\ldots,T_{l}\}$ to be the empty set for all $V_{i}$, $i\in[l]$
3:for every $V_{i}\in Under$ do
4:     $V_{i}\leftarrow V_{i}\cup Vlabel(V_{i})$
5:     if $Tlabel(V_{i})=\emptyset$ then
6:         Let $T^{\prime}_{j}$ be a tree such that $T^{\prime}_{j}\not\in\cup_{s=1}^{l}Tlabel(V_{s})$
7:         $Tlabel(V_{i})\leftarrow T_{j}^{\prime}$
8:for every tree $T_{i}^{\prime}\in\{T_{1}^{\prime},\ldots,T_{l-1}^{\prime}\}$ do
9:     Let $V_{j}$ be the vertex set such that $Tlabel(V_{j})=T_{i}^{\prime}$ (if such a set exists).
10:     if $V_{j}\cap T^{\prime}_{i}=\emptyset$ then
11:         Let $u$ be a vertex of $T_{i}^{\prime}$ such that $u\in N(V_{j}\cap T)$ and $V_{s}$ the vertex set such that $u\in Vlabel(V_{s})$.
12:         if $V_{s}\in Over$ then
13:              $V_{j}\leftarrow V_{j}\cup u$, $Vlabel(V_{s})\leftarrow Vlabel(V_{s})\setminus u$.
14:              if $|V_{j}|=n_{j}$ then
15:                  $Full\leftarrow Full\cup V_{j}$
16:                  if $|Full|$ is equal to the amount of trees the sets in $Full$ intersect then
17:                       Output $Full$
18:                       Run Algorithm 7 with input $(G\setminus Full,C\setminus C^{\prime},N\setminus N^{\prime},\mathcal{T}\setminus\mathcal{T}^{\prime})$ where $C^{\prime}$ is the set of terminals in sets in $Full$, $N^{\prime}$ the size demands of the sets in $Full$ and $\mathcal{T}^{\prime}$ the trees that intersect the sets in $Full$.
19:              if $n_{s}-|V_{s}|>|Vlabel(V_{s})|$ then
20:                  $Over\leftarrow Over\setminus V_{s}$, $Under\leftarrow Under\cup V_{s}$
21:                  Go to line  3
22:         else
23:              $V_{s}\leftarrow V_{s}\setminus u$, $Vlabel(V_{s})\leftarrow Vlabel(V_{s})\setminus u$, $V_{j}\leftarrow V_{j}\cup u$
24:              if $V_{s}\in Full$ then
25:                  $Full\leftarrow Full\setminus V_{s}$                             return $\{V_{1},\ldots,V_{k},Under,Over,Vlabel,Full\}$
Lemma 3.11.
Let a graph $G$, $l$ natural numbers $N=\{n_{1},\ldots,n_{l}\}$, $l$ connected vertex sets $\{V_{1},\ldots,V_{l}\}$, a tree $T$ dominating to $G$ and $l-1$ dominating trees $\{T_{1}^{\prime},\ldots,T_{l-1}^{\prime}\}$ and a function $Vlabel:\{V_{1},\ldots,V_{k}\}\to P(V(G))$, are given as an input to Algorithm 5. Then, after the loop in line 2 (Step 7), each vertex set is connected and is either full or completely contains a dominating tree or is assigned through the function $Vlabel$ at least as many vertices as its remaining size demand that do not belong in any other vertex set.
Proof 3.12.
In order to see that the loop in line 2 terminates, it suffices to notice that it only considers sets in $Under$ and the value of $\sum_{i=1}^{l-1}|T_{i}^{\prime}\cap Tlabel^{-1}(T_{i}^{\prime})|$.
We start by considering the vertex set $V_{j}$ such that $Tlabel(V_{j})=T_{i}^{\prime}$ for some value of $i$. Due to Lemma 3.9, we know that $T_{i}^{\prime}$ is either completely contained in $V_{j}$, in which case we do nothing and we proceed to consider the next tree, or it has a vertex that is adjacent to the part of $V_{j}\cap T_{i}^{\prime}\cup T$ and not in $V_{j}$. We then add that vertex to $V_{j}$ in the same manner as in the loop in line 8 of Algorithm 4 depending on whether this vertex belongs already in some set or not.
Notice that this action preserves the connectivity of the set $V_{j}$ due to the way the vertex is chosen and also does not affect the connectivity of other sets as this is based only on the adjacencies of the parts of each vertex set intersecting $T$, which is not affected. Moreover it increases the value of $\sum_{i=1}^{l-1}|T_{i}^{\prime}\cap Tlabel^{-1}(T_{i}^{\prime})|$ by $1$, which ensures that this loop terminates. Again, every time we add a vertex to a vertex set, it is implied that we also check if the condition to restart Algorithm 7 is met.
Algorithm 5 Sets in Under (Step 6)
1:A $k$-connected graph $G$, a set of $l$ sets $\{T,T^{\prime}_{1},\ldots,T^{\prime}_{l}\}$, $l$ vertices of $G$ $C=\{c_{1},\ldots,c_{l}\}$, $l$ connected vertex sets $\{V_{1},\ldots,V_{l}\}$ such that $c_{i}\in V_{i}$ for all $i\in[l]$ separated in two sets $Under$, $Over$, a function $Vlabel:\{V_{1},\ldots,V_{k}\}\to P(V(G))$, a function $Tlabel:\{V_{1},\ldots,V_{k}\}\to\{T,T^{\prime}_{1},\ldots,T^{\prime}_{l-1}\}$, a set of sets $Full$, a set of $l$ natural numbers $N=\{n_{1},\ldots,n_{l}\}$
2:while $\exists V_{j}\in Under$ such that $V_{j}\not\in Full$ and $V_{j}$ doesn’t completely contain some tree from $\{T_{1}^{\prime},\ldots,T_{l-1}^{\prime}\}$ do
3:     Let $T_{i}^{\prime}$ be the tree such that $Tlabel(V_{j})=T_{i}^{\prime}$
4:     while $V_{j}$ doesn’t contain vertices only from $T_{i}^{\prime}$ and $T_{i}^{\prime}\not\subseteq V_{j}$ and $V_{j}\not\in Full$ do
5:         Let $v\in T_{i}^{\prime}$ such that $v\in N(V_{j}\cap(T\cup T_{i}^{\prime}))$
6:         if $v$ is assigned through the function $Vlabel$ to some set $V_{s}\in Over$ then
7:              $V_{j}\leftarrow V_{j}\cup v$, $Vlabel(V_{s})\leftarrow Vlabel(V_{s})\setminus v$
8:              if $|V_{j}|=n_{j}$ then
9:                  $Full\leftarrow Full\cup V_{j}$
10:                  if $|Full|$ is equal to the amount of trees the sets in $Full$ intersect then
11:                       Output $Full$
12:                       Run Algorithm 7 with input $(G\setminus Full,C\setminus C^{\prime},N\setminus N^{\prime},\mathcal{T}\setminus\mathcal{T}^{\prime})$ where $C^{\prime}$ is the set of terminals in sets in $Full$, $N^{\prime}$ the size demands of the sets in $Full$ and $\mathcal{T}^{\prime}$ the trees that intersect the sets in $Full$.
13:              if $n_{s}-|V_{s}|>|Vlabel(V_{s})|$ then
14:                  $Over\leftarrow Over\setminus V_{s}$, $Under\leftarrow Under\cup V_{s}$
15:                  Run Algorithm 4 with input $(G,\{T,T^{\prime}_{1},\ldots,T^{\prime}_{l-1}\},C,Under,Over,Vlabel,N)$
16:         else
17:              $V_{s}\leftarrow V_{s}\setminus v$, $Vlabel(V_{s})\leftarrow Vlabel(V_{s})\setminus v$, $V_{j}\leftarrow V_{j}\cup v$
18:              if $|V_{j}|=n_{j}$ then
19:                  $Full\leftarrow Full\cup V$
20:                  if $|Full|$ is equal to the amount of trees the sets in $Full$ intersect then
21:                       Output $Full$
22:                       Run Algorithm 7 with input $(G\setminus Full,C\setminus C^{\prime},N\setminus N^{\prime},\mathcal{T}\setminus\mathcal{T}^{\prime})$ where $C^{\prime}$ is the set of terminals in sets in $Full$, $N^{\prime}$ the size demands of the sets in $Full$ and $\mathcal{T}^{\prime}$ the trees that intersect the sets in $Full$.
23:              if $V_{s}\in Full$ then
24:                  $Full\leftarrow Full\setminus V_{s}$                             return $\{V_{1},\ldots,V_{l}\},Over,Under,Full$
Algorithm 6 Remaining Vertices (Step 7)
1:A $k$-connected graph $G$, a set of $l$ dominating trees $\{T,T^{\prime}_{1},\ldots,T^{\prime}_{l}\}$, a set of $l$ vertices $C=\{c_{1},\ldots,c_{l}\}$, $l$ connected vertex sets $\{V_{1},\ldots,V_{l}\}$ such that $c_{i}\in V_{i}$ for all $i\in[l]$ separated to two sets $Under$ and $Over$, a function $Vlabel:\{V_{1},\ldots,V_{k}\}\to P(V(G))$, a function $Tlabel:\{V_{1},\ldots,V_{k}\}\to\mathcal{T}$, a set of vertex sets $Full$, a set of $l$ natural numbers $N=\{n_{1},\ldots,n_{l}\}$
2:for every $V\in Over$ and $V\not\in Full$ do
3:     Let $u$ be a vertex in $Vlabel(V)$ and not in $V$
4:     $V\leftarrow V\cup u$
5:     if $|V|=n_{V}$ then
6:         $Full\leftarrow Full\cup V$
7:         if $|Full|$ is equal to the amount of trees the sets in $Full$ intersect then
8:              Output $Full$
9:              Run Algorithm 7 with input $(G\setminus Full,C\setminus C^{\prime},N\setminus N^{\prime},\mathcal{T}\setminus\mathcal{T}^{\prime})$ where $C^{\prime}$ is the set of terminals in sets in $Full$, $N^{\prime}$ the size demands of the sets in $Full$ and $\mathcal{T}^{\prime}$ the trees that intersect the sets in $Full$.
10:for every $v\in\cup_{i=1}^{l-1}T^{\prime}_{i}\setminus\cup_{i=1}^{l}V_{i}$ do
11:     Let $V$ be a set such that $V\not\in Full$ and $v\in N(V)$
12:     $V\leftarrow V\cup V$
13:     if $|V|=n_{V}$ then
14:         $Full\leftarrow Full\cup V$
15:         if $|Full|$ is equal to the amount of trees the sets in $Full$ intersect then
16:              Output $Full$
17:              Run Algorithm 7 with input $(G\setminus Full,C\setminus C^{\prime},N\setminus N^{\prime},\mathcal{T}\setminus\mathcal{T}^{\prime})$ where $C^{\prime}$ is the set of terminals in sets in $Full$, $N^{\prime}$ the size demands of the sets in $Full$ and $\mathcal{T}^{\prime}$ the trees that intersect the sets in $Full$.               return $\{V_{1},\ldots,V_{l}\}$
Algorithm 7 Gyori-Lovasz theorem for graphs with $k$ dominating trees
1:A $k$-connected graph $G$, $k$ distinct vertices $C=\{c_{1},\ldots,c_{k}\}\in V(G)$, $k$ natural numbers $N=\{n_{1},\ldots,n_{k}\}$ such that $\sum_{i=1}^{k}n_{i}=|V(G)|$, and a set of $k$ vertex disjoint dominating trees of $G$, $\mathcal{T}=\{T_{1},\ldots,T_{k}\}$
2:Initialize the sets $\mathcal{T}_{>1},\mathcal{T}_{1},\mathcal{T}_{0}\leftarrow\emptyset$, $U\leftarrow V(G)\setminus\mathcal{V}(T)$, $Full\leftarrow\emptyset$
[…]
4:$\{V_{1},\ldots,V_{k}\}\leftarrow$ Add Vertices$(\mathcal{T}_{0},\mathcal{T}_{1},\mathcal{T}_{>1},C,\{V_{1},\ldots,V_{k}\},N,U)$
5:($(\{T,T^{\prime}_{1},\ldots,T^{\prime}_{l-1}\},Over,Under,Vlabel)\leftarrow$Choose a Tree Set$(G,\mathcal{T}_{0},\mathcal{T}_{1},\mathcal{T}_{>1},C,N,\{V_{1},\ldots,V_{k}\})$
6:$\{V_{1},\ldots,V_{k},Under,Over,Vlabel,Full\}\leftarrow$Assign Trees$(G,\{T,T^{\prime}_{1},\ldots,T^{\prime}_{l-1}\},C^{\prime}=\{c_{1},\ldots,c_{l}\},Under,Over,Vlabel,N^{\prime}=\{n_{1},\ldots,n_{l}\})$
7:$(V_{1},\ldots,V_{k},Over,Under,Full)\leftarrow$ Sets in Under$(G,\{T,T^{\prime}_{1},\ldots,T^{\prime}_{l-1}\},C^{\prime}=\{c_{1},\ldots,c_{l}\},Under,Over,Vlabel,Tlabel,Full,N^{\prime}=\{n_{1},\ldots,n_{l}\})$
8:$(V_{1},\ldots,V_{l})\leftarrow$Remaining Vertices$(G,\{T,T^{\prime}_{1},\ldots,T^{\prime}_{l-1}\},C^{\prime}=\{c_{1},\ldots,c_{l}\},Under,Over,Vlabel,Tlabel,Full,N^{\prime}=\{n_{1},\ldots,n_{l}\})$
9:Output $V_{1},\ldots,V_{l}$ and restart the algorithm for $G-\cup_{i=1}^{l}V_{i}$ and $k-l$ (if $k-l\neq 0$).
It is important to note that due to the size demands of the sets summing up to less than the sum of the vertices of the participating trees $\{T,T^{\prime}_{1},\ldots,T^{\prime}_{l-1}\}$ there is always a vertex set in $Over$. We are now ready to assemble everything and prove the correctness of Algorithm 7.
Lemma 3.13.
Let a $k$-connected graph with $k$ terminals $C=\{c_{1},c_{2},\ldots,c_{k}\}$, $k$ natural numbers $N=\{n_{1},n_{2},\ldots,n_{k}\}$ such that $\sum_{i=1}^{k}n_{i}=|V(G)|$, and a set of $k$ vertex disjoint dominating trees of $G$, $\mathcal{T}=\{T_{1},T_{2},\ldots,T_{k}\}$ are given as an input of Algorithm 7. It returns a GL partition $V_{1},V_{2},\ldots,V_{k}$ of the graph $G$.
Proof 3.14.
Due to Lemma 3.1 after the first step of our algorithm, each vertex set consists only of its corresponding terminal vertex, and each terminal vertex belongs in some tree. The second step of the algorithm adds all vertices of trees containing at least one terminal to some set while preserving the connectivity of the vertex sets due to Lemma 3.3 as the input sets contain only the corresponding terminal and are hence connected. Moreover, the same holds for the addition of the vertices not in any tree on those sets (both in preserving connectivity and in running time) due to Lemma 3.5.
The next routine our algorithm uses is Algorithm 3. The trees are partitioned into at most $k/2$ sets, and after choosing at least one, we consider every vertex of the trees contained in this set. Due to Lemma 3.7, we know that the while loop of this algorithm terminates after at most $n$ iterations. Moreover, since the input of this routine is as desired, due to Lemma 3.5, the function $Vlabel$ results in at least one connected vertex set being in $Over$ while also the vertex sets remain connected if all of the assigned vertices were added to them.
In Algorithm 4, if some set moves from $Over$ to $Under$, then algorithm 3 is called, which happens at most $k$ times. As stated in Lemma 3.9, since the input of this algorithm is the desired one, this results in a function $Tlabel$ with the property that each vertex set is either in $Over$ or assigned a tree through this function and it also contains a vertex from that tree.
Afterward, by running Algorithm 5, each vertex set considered is connected and either full or completely contains a dominating tree or is in $Over$.
The rest of the algorithm (Algorithm 6, Step 7) is relatively straightforward as first we add to each vertex set in $Over$ as many vertices as needed for it to meet its size demand and then due to Lemma 3.11 we know that each remaining non-full set contains a dominating tree. Hence, we can arbitrarily add vertices to the non-full sets until they meet their size demand while preserving the connectivity of each vertex set.
Due to our tree set satisfying the condition of line 8, we know that all of them are full at the end of this procedure. Hence, we have $l$ full sets intersecting exactly $l$ dominating trees, which is the desired condition to output the so far full sets and restart the algorithm for the altered values of $k$, $G$, $\mathcal{T}$, $C$ and $N$. After at most $k$ times repetition of Algorithm 7, we get the required partition $V_{1},V_{2},\ldots,V_{k}$.
We conclyde this section by showing that the Algorithm 7 runs in polynomial time.
Running Time of Algorithm 7: Note that Algorithm 1 runs in time $\mathcal{O}(nk)$ as each tree is considered once and there are at most $k$ terminals not already in some tree. Next, one application of Algorithm 2 runs in $\mathcal{O}(nk)$ time because each vertex of the dominating trees is considered and adds vertices from its neighborhood in the tree. Next, Algorithm 3 finds a set of trees from $\mathcal{T}_{0}$ corresponding to each tree in $T_{>1}$, partitioning eventually the given set of dominating trees.Therefore, it takes overall time $\mathcal{O}(nk)$.
One application of Algorithm 4 runs in $\mathcal{O}(k^{2}n^{2})$ time as each tree of the chosen set is considered and exactly one vertex is considered and then added to it. The running time of Algorithm 5 is $\mathcal{O}(n^{3})$ as one can notice from the termination condition equivalent to the while loop of the algorithm described in the proof of Lemma 3.11.
Observe that any algorithm calls Algorithm 7 when any vertex set get full, that means, at most $k$ times. It follows that any algorithm as part of the Algorithm 7 called at most $k$ times. Thus, the total running time is $\mathcal{O}(k(k^{2}n^{2}+n^{3}))$, that is, $\mathcal{O}(n^{5})$.
4 Győri–Lovász Partition for Interval Graphs
In this section, we show that every $k$-connected interval graph admits a CDS partition of size $k$ (Theorem 1.6), which leads to an efficient construction for the Győri–Lovász theorem on this graph class (Theorem 1.10). We now restate the main result of this section for convenience.
See 1.6
Algorithm 8 $k$-CDS Interval Graph
1:A $k$-connected interval graph $G$
2:Compute a minimum width path decomposition $B_{1},\dots,B_{\ell}$ in $G$.
3:$V(G)\leftarrow V(G)\cup\{s,t\}$, $E(G)\leftarrow E(G)\cup\{sv:\text{ for all }v\in B_{1}\}\cup\{tv:\text{ for all }v\in B_{\ell}\}$
4:Compute $k$ disjoint $s-t$ paths in $G$, $\mathcal{P}=P_{1},\dots,P_{k}$ and delete the endpoints $s$ and $t$ from each path.
5:return $\mathcal{P}$
The algorithm takes advantage of a path decomposition of the graph. To prove Theorem 1.6, we show that the Algorithm 8 produces $k$ dominating paths in polynomial time.
Lemma 4.1.
Algorithm 8 returns $k$ disjoint paths, each dominating $V(G)$, and runs in polynomial time.
Proof 4.2.
It is well-known that in a minimum width path decomposition of an interval graph $G$, the induced subgraph of each bag $B_{i}$ forms a clique in $G$ [20]. Furthermore, each bag acts as a separator in $G$, meaning that a path from $s$ to $t$ must cross every bag. Combining both facts, we conclude that a path $P$ from $s$ to $t$, excluding the endpoints $s$ and $t$, dominates $V$.
Since every bag $B_{i}$ is a separator, and by the $k$-connectivity of $G$, each bag has size at least $k$. Adding a vertex to a $k$-connected graph and connecting it to at least $k$ vertices that preserves $k$-connectivity. As a result, there are $k$ disjoint paths between $s$ and $t$.
Computing a minimum-width path decomposition for interval graphs can be done in polynomial time [20]. Computing $k$ disjoint paths can also be done in polynomial time by using a maximum flow algorithm. Thus, the algorithm runs in polynomial time.
5 Győri–Lovász Partition for Biconvex Graphs
This section is dedicated to the result for biconvex bipartite graphs. We show that every $k$-connected biconvex bipartite graph admits a CDS partition of size $k$ (Theorem 1.5), which leads to an efficient construction for the Győri–Lovász theorem on this graph class (Corollary 1.7). We now restate the main result of this section for convenience.
See 1.5
Algorithm 9 $k$-CDS Biconvex
1:A $k$-connected biconvex graph $G=(A\cup B,E)$ with biconvex ordering $\sigma$ such that $\sigma_{A}=(a_{1},\dots,a_{n})$ and $\sigma_{B}=(b_{1},\dots,b_{m})$
2:Compute $k$ disjoint $a_{1}-a_{n}$ induced paths $\mathcal{T}=T_{1},\dots,T_{k}$ and delete the endpoints $a_{1}$ and $a_{n}$ from each path
3:for every $T_{i}\in\mathcal{T}$ do
4:     if $T_{i}\cap N(b_{1})=\emptyset$ then
5:         $T_{i}\leftarrow T_{i}\cup v$ such that $v\in N(b_{1})$ and $v\not\in V(\mathcal{T})$
6:     if $T_{i}\cap N(b_{m})=\emptyset$ then
7:         $T_{i}\leftarrow T_{i}\cup v$ such that $v\in N(b_{m})$ and $v\not\in V(\mathcal{T})$
8:return $\mathcal{T}$
To prove the Theorem 1.5, we design an algorithm to construct a CDS partition for biconvex graphs, given in Algorithm 9. To prove the correctness of Algorithm 9, first we need to show that there always exists a neighbor of $b_{1}$ or $b_{m}$ in $V(G)\setminus V(\mathcal{T})$ to add, at the lines 4 and 6.
Lemma 5.1.
Let $G$ be a connected biconvex graph $G=(A\cup B,E)$ and $\sigma$ be a biconvex ordering such that $\sigma_{A}=(a_{1},\dots,a_{n})$ and $\sigma_{B}=(b_{1},\dots,b_{m})$. Let $P^{\prime}$ be an induced $(a_{1},a_{n})$-path and $P=P^{\prime}\setminus\{a_{1},a_{n}\}$. Then the path $P$ contains at most one neighbor of $b_{1}$ and at most one neighbor of $b_{m}$. Furthermore, the path $P$ dominates the vertex set $A$.
Proof 5.2.
Let $P$ be such a path and assume that $|V(P)\cap N(b_{1})|>1$. Let $a^{\prime}_{1},a^{\prime}_{2}\in N(b_{1})\cap V(P)$. Since $G$ is bipartite, $a^{\prime}_{1},a^{\prime}_{2}\in A$. Let $b^{\prime}_{1},b^{\prime}_{2},b^{\prime}_{3},b^{\prime}_{4}\in B$ be the neighbors of $a^{\prime}_{1},a^{\prime}_{2}$ in $P$, such that $b^{\prime}_{1}\leq b^{\prime}_{2}\leq b^{\prime}_{3}\leq b^{\prime}_{4}$. Note that the $b^{\prime}_{i}$’s are not necessarily distinct; however, at most one of them has $a^{\prime}_{1}$ and $a^{\prime}_{2}$ as neighbors in $P$. Without loss of generality, let $a^{\prime}_{1}$ be the neighbor of $b^{\prime}_{4}=b_{x}$. Consequently, the neighborhood of $a^{\prime}_{1}$ in $G$ contains $b_{1},\dots,b_{x}$ by convexity, and in particular, the vertices $b^{\prime}_{1},b^{\prime}_{2},b^{\prime}_{3},b^{\prime}_{4}$. Thus, the degree of $a^{\prime}_{1}$ is at least three in $G[V(P)]$, contradicting the precondition that $P$ is an induced path.
Symmetrically, we can make the same argument for $b_{m}$. Suppose $|V(P)\cap N(b_{m})|>1$. Let $a^{\prime\prime}_{1},a^{\prime\prime}_{2}\in N(b_{m})$ be two vertices in $P$. Since $G$ is bipartite, $a^{\prime\prime}_{1},a^{\prime\prime}_{2}\in A$. Let $b^{\prime\prime}_{1},b^{\prime\prime}_{2},b^{\prime\prime}_{3},b^{\prime\prime}_{4}\in B$ be the neighbors of $a^{\prime\prime}_{1},a^{\prime\prime}_{2}$ in $P$, such that $b^{\prime\prime}_{1}\leq b^{\prime\prime}_{2}\leq b^{\prime\prime}_{3}\leq b^{\prime\prime}_{4}$. Note that the $b^{\prime\prime}_{i}$’s are not necessarily distinct; however, at most one of them has $a^{\prime\prime}_{1}$ and $a^{\prime\prime}_{2}$ as neighbors in $P$. Without loss of generality, let $a^{\prime\prime}_{1}$ be the neighbor of $b^{\prime\prime}_{1}=b_{y}$. Consequently, the neighborhood of $a^{\prime\prime}_{1}$ in $G$ contains $b_{y},\dots,b_{m}$ by convexity, and in particular, the vertices $b^{\prime\prime}_{1},b^{\prime\prime}_{2},b^{\prime\prime}_{3},b^{\prime\prime}_{4}$. Thus, the degree of $a^{\prime\prime}_{1}$ is at least three in $G[V(P)]$, contradicting the precondition that $P$ is an induced path.
Next, we want to show that the path $P$ dominates the vertex set $A$. Clearly, $a_{1}$ and $a_{n}$ are dominated by $P$. Assume there exists a vertex $a_{i}\in A\setminus\{a_{1},a_{n}\}$ that is not dominated by $P$. Originally, the path extended from $a_{1}$ to $a_{n}$, which implies the existence of consecutive vertices $a_{x},b_{y},a_{z}$ in the original $P$ with $a_{x}<a_{z}$ such that $a_{x}\leq a_{i}\leq a_{z}$. However, by convexity $a_{i}\in N(b_{y})$, contradicting our assumption.
We now use the fact that $k$-connectivity ensures each vertex has at least $k$ neighbors. By Lemma 5.1, each $T\in\mathcal{T}$ contains at most one neighbor of $b_{1}$ or $b_{m}$. Consequently, if $T$ has no neighbor of $b_{1}$ or $b_{m}$, there must be an available neighbor in $A$ in step 3 as $|\mathcal{T}|=k$.
The sets in $\mathcal{T}$ are disjoint by construction. Recall that every $k$-connected graph guarantees the existence of at least $k$ disjoint paths between any pair of vertices. Since every path $T$ computed in step 2 is connected and already dominates $A$ (see Lemma 5.1), every additional vertex in $A$ that is added to $T$ preserves its connectivity. Next, we show that each set in $\mathcal{T}$ dominates the whole graph.
Lemma 5.3.
Let $\mathcal{T}$ be a returned set by Algorithm 9. Then, every vertex set $T\in\mathcal{T}$ dominates the input graph $G$.
Proof 5.4.
Let $T\in\mathcal{T}$. By Lemma 5.1, $T$ dominates $A$ as $T$ contains an $(a_{1},a_{n})$-path except $\{a_{1},a_{n}\}$. Now, it remains to show that $T$ dominates $B$. Let $b_{x}$ and $b_{y}$ be the vertices with the lowest and highest ordering value in $V(T)\cap B$. First, consider the case $b_{x}=b_{y}$. By the construction of $T$ in the algorithm, $N(V(T))$ contains $b_{1}$ and $b_{m}$, and by Lemma 5.1, $b_{x}$ dominates $A$. Consequently, $T$ dominates $B$, since there exists a path $b_{1},a,b_{x}$ and a path $b_{x},a^{\prime},b_{m}$, where $a,a^{\prime}\in V(T)$.
Now, consider the case $b_{x}<b_{y}$. We begin by proving that the sequence $b_{x},\dots,b_{y}$ is dominated by $T$. Note that $b_{x}$ and $b_{y}$ are in $T$ before loop at line 3, since in this step only vertices from $A$ are added. That is, $b_{x}$ and $b_{y}$ are part of the path computed in step 2, and therefore, there exists a path $P$ in $T$ from $b_{x}$ to $b_{y}$. Suppose there exists a vertex $b_{i}\in\{b_{x},\dots,b_{y}\}$ that is not dominated by $P$. Since there is a path from $b_{x}$ to $b_{y}$, there exist consecutive vertices $b_{j},a_{z},b_{k}$ in $P$ with $b_{j}<b_{k}$ such that $b_{j}\leq b_{i}\leq b_{k}$. However, by convexity $b_{i}\in N(a_{z})$, contradicting our assumption. Finally, by construction $N(V(T))$ contains $b_{1}$ and $b_{m}$, and by Lemma 5.1, $T$ dominates $A$. Consequently, $T$ dominates $B$, since there exists a path $b_{1},a,b$ and a path $b^{\prime},a^{\prime},b_{m}$, where $a,a^{\prime},b,b^{\prime}\in V(T)$.
By using a maximum flow algorithm, we can compute $k$ disjoint paths between $a_{1}$ and $a_{n}$ in polynomial time. In the final step, we can check if a $T\in\mathcal{T}$ has no neighbor in $b_{1}$ or $b_{m}$ and add extra vertex in $T$, in polynomial time. Hence, Algorithm 9 runs in polynomial time. The collection returned by Algorithm 9, consists of $k$ disjoint connected dominating sets, which can be extended to a CDS partition of $G$. Hence, Theorem 1.5 follows.
6 Conclusion
As we showcase in this work, the algorithm of Theorem 1.3 opens up a new avenue for efficiently computing connected partitions, in the settings where sufficiently large CDS partitions are available. This motivates further research into improving bounds for CDS partitions beyond the worst-case instances, as well as into parameterized and approximation algorithms for computing a CDS partition of the largest size.
To the best of our knowledge, Corollaries 1.4 and 1.10 are the first examples of a Győri–Lovász type of results that allow for a trade-off between efficiency of the construction and the connectivity requirement. It would be interesting to investigate whether a more refined trade-off is possible, interpolating smoothly between the polynomial time for higher connectivity and exponential time for extremal connectivity.
Our work also initiates the study of the gap between connectivity and the size of a CDS packing in structured graph classes. It remains a curious open question to characterize the class of graphs, for which $k$-connectivity implies the existence of a $k$-CDS partition. So far, our results show that this class contains interval and biconvex graphs (Theorems 1.6 and 1.5), but does not contain chordal and convex bipartite graphs.